a simple pendulum is set upon trolleywhich slides down a frictionless inclined plane making an angle @ with the horizontal.then the time period of pendulum is ?

Grade:

2 Answers

Chetan Mandayam Nayakar
11 years ago

The correct answer is: 

Answer=2(pi)sqrt(l/G),(G)2=(g-g(sin2(@)))2+(g(sin@(cos@)))2




ruja cool
11 years ago